tdaniels0974:

Renata wins a $20 gift card to an online music site. After Renata purchases 16 songs, the gift card has a remaining balance of $0. Which equation represents the relationship between y, the remaining balance on Renata’s gift card, and x, the number of songs purchased?

Leah999:

.__________________________. idek

XioGonz:

Renata wins a $20 gift card After purchasing 16 songs, there is a remaining balance of $0 So $20/16 =$5/4 =$1.25 2s=$1.25 3s=$1.25 Now since Renata had $20, to begin with: Her outcome would be 5x + 4y = 80